Grade 11Physical Chemistry

When hno3 is converted into nh3, the equivalent weight of hno3 will be? (m=molecular weight of hno3)? Plzz explain the steps

Spandan
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
To calculate equivalent weight We need to calculate the Valence factor first for the N atom.
Therfore the oxidation state for N in reactant side is => +5
In the product side the oxidation state is => -3
Therfore gthe valence factor would be = > (5+3) = 8
The Eq. weight => m/8
Where m is the molecular wieght given.
Nehal  Gangwal
7 Years ago
The equivalent weight of HNO3= Molar mass/valency factor
Valency factor of HNO3= the O.N. of Nitrogen on both sides= 5+3= 8
Therefore, the equivalent weight of HNO3= Molar mass (m) / 8